A large number of applicants for admission to graduate study in business are given an aptitude test. Scores are normally distributed with a mean of 460 and standard deviation of 80. What fraction of the applicants would you expect to have a score of 400 or above
Answer #1

Solution :

Given that ,

mean = \mu = 460

standard deviation = \sigma = 80

P(x \geq 400) = 1 - P(x  \leq 400)

= 1 - P[(x - \mu ) / \sigma \leq (400 - 460) / 80]

= 1 -  P(z \leq -0.75)   

= 1 - 0.2266

= 0.7734

Answer = 0.7734